The length of a new rectangular playing field is 7yards longer than quadruple the width. If the perimeter of the rectangular playing field is 564 ​yards, what are its​ dimensions?
Answer by ikleyn(52776)

.
The length of a new rectangular playing field is 7 yards longer than quadruple the width. If the perimeter of the rectangular
playing field is 564 ​yards, what are its​ dimensions?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let L be the length, and W be the width.

The condition says 

L = 4W + 7.

Then the perimeter is

2L + 2W = 2*(4W+7) + 2W = 8W + 14 + 2W = 10W + 14.

And the equation is 

10W + 14 = 564,  or

10W = 564 - 14,

10W = 550, 

W = 55 yards.

Now you find L.